OpenStudy (livya15):

I need help refreshing my memory on how to simplify these fractions with a rational exponent in denominator. Will give MEDAL!

OpenStudy (livya15):

\[\frac{ 1 }{ t ^{5/2} }\]

OpenStudy (livya15):

\[\frac{ 1 }{ 2t ^{5/2} }\]

jimthompson5910 (jim_thompson5910):

do they want you to convert to radical form?

OpenStudy (livya15):

well, for the second fraction the answer is: \[\frac{ 1 }{ t ^{2}\sqrt{t} }\] I want to know how to get to that point.

jimthompson5910 (jim_thompson5910):

oh I gotcha

jimthompson5910 (jim_thompson5910):

rule: \[\LARGE x^{m/n} = \sqrt[n]{x^m}\] in this case, m = 5 and n = 2 \[\LARGE x^{5/2} = \sqrt[2]{x^5}\] \[\LARGE x^{5/2} = \sqrt{x^5}\]

jimthompson5910 (jim_thompson5910):

then you simplify the radical \[\Large \sqrt{x^5} = \sqrt{x^4*x}\] \[\Large \sqrt{x^5} = \sqrt{x^4}*\sqrt{x}\] \[\Large \sqrt{x^5} = x^2\sqrt{x}\]

OpenStudy (livya15):

that makes sense to me now! thank you!

jimthompson5910 (jim_thompson5910):

do you see what to do next?

jimthompson5910 (jim_thompson5910):

well I guess it depends on the book, so it's optional, but some books rationalize the denominator to do that, you'd multiply top and bottom by sqrt(t) \[\Large \frac{ 1 }{ t ^{2}\sqrt{t} } = \frac{ 1*\sqrt{t} }{ t ^{2}\sqrt{t}*\sqrt{t} }\] \[\Large \frac{ 1 }{ t ^{2}\sqrt{t} } = \frac{ \sqrt{t} }{ t ^{2}*t }\] \[\Large \frac{ 1 }{ t ^{2}\sqrt{t} } = \frac{ \sqrt{t} }{ t ^{3} }\]